# Subspace-Protected Topological Phases and Bulk-Boundary Correspondence

### v1

@@ Abstract
While tremendous research has revealed that symmetry enriches topological phases of matter, more general principles that protect topological phases have yet to be explored. In this Letter, we elucidate the roles of subspaces in free-fermionic topological phases. A subspace property for Hamiltonians enables us to define new topological invariants. They result in peculiar topological boundary phenomena, i.e., the emergence of an unpaired zero mode or zero-winding skin modes, characterizing subspace-protected topological phases. We establish and demonstrate the bulk-boundary correspondence in subspace-protected topological phases. We further discuss the interplay of the subspace property and internal symmetries. Toward application, we also propose possible platforms possessing the subspace property.
Symmetry is a fundamental principle to characterize and classify topological phases of matter [1]. In particular, various types of symmetry provide free-fermion systems with rich topological structure, as in topological insulators and superconductors [2, 3, 4, 5, 6]. A unitary symmetry resolves the Hilbert space into sectors, in which topological phases are individually classified on the basis of internal symmetries or crystalline symmetries [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. A significant characteristic of topological phases is the bulk-boundary correspondence: the nontrivial topological phases in gapped bulk involves the presence of gapless boundary modes [20, 21, 22, 23, 24, 25, 26, 27]. Nontrivial topological insulators protected by time-reversal symmetry (TRS) exhibit helical gapless modes as the Kramers pairs [28, 29, 30, 31, 32, 33, 34, 35, 36, 37], and nontrivial topological superconductors can host Majorana zero modes reflecting particle-hole symmetry [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50].
@@
However, several topological phenomena in free-fermion systems challenge explanation beyond conventional notion of symmetry. For instance, recent studies have observed topological zero modes protected by a so-called sub-symmetry in some Hermitian systems with broken sublattice symmetry [113, 114, 115, 116, 116]. In non-Hermitian cases, trivial point-gapped systems have been reported to host skin modes even with no symmetry [117, 118]. Thus, exploring general principles protecting topological phenomena is an elusive but urgent issue.
In this Letter, we reveal that subspaces of Hilbert spaces serve as a principle to introduce a novel type of free-fermion topological phases in both Hermitian and non-Hermitian systems beyond conventional notion of symmetry. Specifically, we consider the following condition, which we call the subspace property, for a Bloch Hamiltonian $H(\boldsymbol{k})$ acting on a Hilbert space $\mathcal{H}$ of internal degrees of freedom,
@@
where $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are subspaces of $\mathcal{H}$ independent of the wave vector $\boldsymbol{k}$, as illustrated in Fig. 1 (a). We propose the concept of topological phases protected by the subspace property (1), dubbed subspace-protected topological phases, and establish the bulk-boundary correspondence in the phases. Furthermore, we clarify the interplay of subspace-protected topological phases and internal symmetries. Toward application, we provide guideline for the search of physical platforms possessing the subspace property.
Figure 1: (a) Schematic figure of the subspace property with two-dimensional subspaces $\mathcal{M},\mathcal{M}^{\prime}$ in the three-dimensional Hilbert space $\mathcal{H}$. (b) Classification of relations between $\mathcal{M}$ and $\mathcal{M}^{\prime}$. In the case of $\dim\mathcal{M}=\dim\mathcal{M}^{\prime}$, the subspace-protected topological invariant is well-defined.
@@
Discussion.— In summary, we have discovered subspace-protected topological phases and established the bulk-boundary correspondence. Our work reveals that a principle beyond symmetry plays an important role in topological phases of free-fermion systems.
@@
Through the Letter, we have assumed in the subspace property (1) that $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are complex subspaces. Taking $\mathcal{M}$ and $\mathcal{M}^{\prime}$ as real or quaternionic subspaces may enrich subspace-protected topological phases. Futhermore, $K$-theoretical classification of the phases is also an open problem.
Note added.—After the completion of this work, we became aware of a recent related work [153].

### v2

@@ Abstract
While tremendous research has revealed that symmetry enriches topological phases of matter, more general principles that protect topological phases have yet to be explored. In this Letter, we elucidate the roles of subspaces in free-fermionic topological phases. A subspace property for Hamiltonians enables us to define new topological invariants. It results in peculiar topological boundary phenomena, i.e., the emergence of an unpaired zero mode or zero-winding skin modes, characterizing subspace-protected topological phases. We establish and demonstrate the bulk-boundary correspondence in subspace-protected topological phases. We further discuss the interplay of the subspace property and internal symmetries. Toward application, we also propose possible platforms possessing the subspace property.
Symmetry is a fundamental principle to characterize and classify topological phases of matter [1]. In particular, various types of symmetry provide free-fermion systems with rich topological structure, as in topological insulators and superconductors [2, 3, 4, 5, 6]. A unitary symmetry resolves the Hilbert space into sectors, in which topological phases are individually classified on the basis of internal symmetries [7, 8, 9, 10, 11, 12] or crystalline symmetries [13, 14, 15, 16, 17, 18, 19, 20, 21]. A significant characteristic of topological phases is the bulk-boundary correspondence: the nontrivial topological phases in gapped bulk involves the presence of gapless boundary modes [22, 23, 24, 25, 26, 27, 28, 29]. Nontrivial topological insulators protected by time-reversal symmetry (TRS) exhibit helical gapless modes as the Kramers pairs [30, 31, 32, 33, 34, 35, 36, 37, 38, 39], and nontrivial topological superconductors can host Majorana zero modes reflecting particle-hole symmetry [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52].
@@
However, several topological phenomena in free-fermion systems challenge explanation beyond conventional notion of symmetry. For instance, recent studies have observed topological zero modes protected by a so-called sub-symmetry in some Hermitian systems with broken sublattice symmetry [115, 116, 117, 118, 119, 120]. In non-Hermitian cases, trivial point-gapped systems have been reported to host skin modes even with no symmetry [121, 122]. Thus, exploring general principles protecting topological phenomena is an elusive but urgent issue.
In this Letter, we reveal that subspaces of Hilbert spaces serve as a principle to introduce a novel type of free-fermion topological phases in both Hermitian and non-Hermitian systems beyond the conventional notion of symmetry. Specifically, we consider the following condition, which we call the subspace property, for a Bloch Hamiltonian $H(\boldsymbol{k})$ acting on a Hilbert space $\mathcal{H}$ of internal degrees of freedom,
@@
where $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are subspaces of $\mathcal{H}$ independent of the wave vector $\boldsymbol{k}$, as illustrated in Fig. 1 (a). Note that the subspace property is not reduced to group-theoretical symmetry operations. We propose the concept of topological phases protected by the subspace property (1), dubbed subspace-protected topological phases, and establish the bulk-boundary correspondence in the phases. Furthermore, we clarify the interplay of subspace-protected topological phases and internal symmetries. Toward application, we provide searching guidelines for physical platforms possessing the subspace property.
Figure 1: (a) Schematic figure of the subspace property with two-dimensional subspaces $\mathcal{M},\mathcal{M}^{\prime}$ in the three-dimensional Hilbert space $\mathcal{H}$. The subspace property is regarded as a geometric condition on the Hilbert space. (b) Classification of relations between $\mathcal{M}$ and $\mathcal{M}^{\prime}$. In the case of $\dim\mathcal{M}=\dim\mathcal{M}^{\prime}$, the subspace-protected topological invariant is well-defined.
@@
Discussion.— In summary, we have discovered subspace-protected topological phases and established the bulk-boundary correspondence. Our work reveals that the new principle beyond symmetry may play an essential role in topological phases.
@@
Through the Letter, we have assumed in the subspace property (1) that $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are complex subspaces. Taking $\mathcal{M}$ and $\mathcal{M}^{\prime}$ as real or quaternionic subspaces may enrich subspace-protected topological phases. Furthermore, $K$-theoretical classification of the phases is also an open problem.
Note added.— A part of this work was presented in [157]. After completing this work, we became aware of a recent work [158]. Although Ref. [158] introduces a similar terminology ”subspace symmetry”, it is an emergent symmetry of a Hamiltonian, not a property of a subspace. Thus, the concept is entirely different from ours.